OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y = x2 - x - 56. A. -56 B. 56 C. -225/4 D. 225/4

OpenStudy (btaylor):

the x-value of the minimum is at -b/2a, so here 1/2. The y-value here is -225/4, or C

OpenStudy (anonymous):

y=x^2-x+1/4-1/4-56 =(x-1/2)^2-(1+224)/4 =(x-1/2)^2-225/4 minimum value of y=-225/4 and is attained at x=1/2